# Proactive Model Adaptation Against Concept Drift for Online Time Series Forecasting

### v2

@@ E.2. Performance with Partial Ground Truth
In this section, we introduce variants that update the model with feedback on $\mathbf{Y}_{t-H}$ and partial ground truth $\{\tilde{\mathbf{Y}}_{t-i}\}_{i=1}^{H-1}$, where ${\tilde{\mathbf{Y}}_{t-i}=[v_{t-i},\cdots,v_{t-1}]}$. Since each update involves $H$ training samples, the GPU memory overhead becomes about $H$ times than one involving $\mathbf{Y}_{t-H}$ only. We leave out results on ECL and Traffic datasets, where the GPU memory overhead exceeds the limit of our Nvidia 4090 GPUs (24GB). As shown in Table 8, Proceed consistently outperforms online gradient descent method under this setting. Compared with Table 3, using both $\mathbf{Y}_{t-H}$ and partial ground truth for feedback-based adaptation can improve the performance, while the improvement is insignificant. This indicates that samples with partial ground truth cannot reveal the concept of the test sample, necessitating proactive model adaptation. Though the partial ground truth can be beneficial to our method, the finetuning cost scales up with $H$ (e.g., up to either 96$\times$ latency or 96$\times$ GPU memory). Thus, it would be much more efficient if using $\mathbf{Y}_{t-H}$ only. Notably, our proposed method using $\mathbf{Y}_{t-H}$ only can still outperform GD using $H$ training samples in most cases. .
E.3. Variants of the Concept Encoder
In this section, we study different kinds of operations on multivariate time series using our concept encoder. Let $\mathbf{c}_{t,i}$ denote the latent concept feature vector of the $i$-th variate. Apart from the average introduced in Eq. (1), we implement two variants as follows:
  • linear transformation: we learn a linear layer that transforms the concatenation of all latent features, i.e., $c_{t}=W^{\top}[c_{t0},c_{t1},\cdots]$.
  • weighted sum: we learn an individual weight $w_{i}\in\mathbb{R}$ for each $i$-th variate, and $c_{t}=\sum_{i}w_{i}c_{ti}$.
In Table 9, we compare the performance of the variants in terms of the average MSE with $H\in$ 24,48,96. The results demonstrate that the average operation is simple yet effective. One possible reason is that all variates are typically assumed to have equal importance during both training and evaluation.
Table 9. Performance comparison between variants of concept encoders..
Model Method ETTh2 ETTm1 Weather ECL Traffic
PatchTST linear trans. 3.498 0.570 0.985 5.661 0.363
weighted sum 3.467 0.571 0.986 5.435 0.362
average 3.468 0.554 0.986 4.732 0.356
iTransfo. linear trans. 4.400 0.548 1.018 4.944 0.355
weighted sum 4.441 0.551 1.021 5.260 0.352
average 4.216 0.543 1.017 4.752 0.352

### v3

@@ E.2. Performance with Partial Ground Truth
In this section, we introduce variants that update the model with feedback on $\mathbf{Y}_{t-H}$ and partial ground truth $\{\tilde{\mathbf{Y}}_{t-i}\}_{i=1}^{H-1}$, where ${\tilde{\mathbf{Y}}_{t-i}=[\mathbf{v}_{t-i},\cdots,\mathbf{v}_{t-1}]}$. Since each update involves $H$ training samples, the GPU memory overhead becomes about $H$ times than one involving $\mathbf{Y}_{t-H}$ only. We leave out results on ECL and Traffic datasets, where the GPU memory overhead exceeds the limit of our Nvidia 4090 GPUs (24GB). As shown in Table 9, Proceed consistently outperforms the online gradient descent method under this setting. Compared with Table 3, using both $\mathbf{Y}_{t-H}$ and partial ground truth for feedback-based adaptation can improve the performance, while the improvement is insignificant. This indicates that samples with partial ground truth cannot reveal the concept of the test sample, necessitating proactive model adaptation. Though the partial ground truth can be beneficial to our method, the finetuning cost scales up with $H$ (e.g., up to either 96$\times$ latency or 96$\times$ GPU memory). Thus, it would be much more efficient if using $\mathbf{Y}_{t-H}$ only. Notably, our proposed method using $\mathbf{Y}_{t-H}$ only can still outperform GD using $H$ training samples in most cases. .